OpenStudy (anonymous):

Solve the equation on the interval 0≤θ<2π 2cos^2 theta + cos theta = 0

OpenStudy (mathstudent55):

\( 2 \cos^2 \theta + \cos \theta = 0\) Treat \( \cos \theta\) as a variable, and solve the quadratic equation for \( \cos \theta \) by factoring. Then use trig to solve for \( \theta\).

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm